Tomas's age is 33 and Tim is 6.In how many years will Tomas's age be 4 times more than Tim's
Answer by mananth(16946)

Tomas's age is 33
Tim is 6.
After x years
Tomas's age = 33+x
Tim's age = 6+x
4(6+x)=33+x
24+4x=33+x
3x=9
x=3 after 3 years